a) Solve the differential equation
dy/dx+(x/1-x^2)y=x (sqrt y),y(0)=1
1
Expert's answer
2020-03-27T14:32:08-0400

Given Differential equation is,


"\frac{dy}{dx} +( \frac {x}{1 -x^2}) y= x \times \sqrt y"

Divide by "\sqrt y" on both sides of the above equation, then we get



"\frac {1}{\sqrt y} \frac{dy}{dx} + (\frac {x}{1 -x^2}) \sqrt y = x .............(A)"

Let "\sqrt y = t"

Differentiate with respect to "x"

Then,

"\frac {1}{2\sqrt y} \frac {dy}{dx} = \frac {dt}{dx}"

Then, the equation (A) becomes



"2 \frac {dt}{dx} + (\frac {x}{1 -x^2}) t = x"

divide by 2


"\frac {dt}{dx} + \frac {1}{2} (\frac {x}{1 -x^2}) t = \frac {x}{2}"

This is the Linear Differential equation in the form "\frac {dt}{dx} + P(x) t= Q(x)"


Integrating factor = I.F = "e^ {\int p(x) dx} = e^ {\int \frac {1}{2} (\frac {x}{1 -x^2})dx}"


"= e^{\frac {1}{4}\int \frac {2x}{1-x^2}dx } = e^ {- \frac {1}{4} ln {( 1- x^2) }} = \frac {1}{\sqrt[4] {1-x^2}}"

Now the solution of the linear differential equation is

"t \times (I.F)= \int (I.F) \times Q \times dx + c"

"t \times \frac {1}{\sqrt[4] {1- x^2}} = \int \frac {1}{\sqrt[4] {1- x^2}} \times \frac {x}{2} \space dx +c"

 



"\frac {t}{\sqrt [4] {1-x^2}} = -\frac {1}{4}\int \frac {-2x}{\sqrt [4]{1-x^2}} dx + c"

use substitution method on right side Integration,

Let "1 - x^2 = m \space then \space - 2x dx = dm"


"\frac {t}{\sqrt [4] {1-x^2}} = -\frac {1}{4}\int \frac {1}{\sqrt [4]{m}} dm + c"

"\frac {t}{\sqrt [4] {1-x^2}} = -\frac {1}{4} \times \frac {m^ {\frac {1}{4} + 1}}{\frac {1}{4}+1}+ c"

"\frac {t}{\sqrt [4] {1-x^2}} = -\frac {4}{4} \times \frac {m^ {\frac {5}{4} }}{5}+ c"

"\frac {t}{\sqrt [4] {1-x^2}} = -\frac {(1-x^2)^ {\frac {5}{4} }}{5}+ c"

"y(0)=1"



"\frac {1}{\sqrt [4] {1-0}} = -\frac {(1-0)^ {\frac {5}{4} }}{5}+ c"

"1 = \frac{-1}{5} + c \\
c = 1 + \frac {1}{5} = \frac {6}{5}"

So the solution of the equation is,


"\frac {\sqrt {y}}{\sqrt [4] {1-x^2}} = -\frac {(1-x^2)^ {\frac {5}{4} }}{5}+ \frac{6}{5}"